Circular No. 7276

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 P. M. Garnavich and K. Z. Stanek, Harvard-Smithsonian Center for

Astrophysics; and P. Berlind, Fred L. Whipple Observatory (FLWO), report the detection of an optical counterpart to the new x-ray transient XTE J1859+226. CCD images taken on Oct. 12.07 UT with the FLWO 1.2-m telescope show a bright stellar object not seen on the second Palomar Sky Survey. The exact position based on the HST Guide Star Catalog is α = 18^{h}58^{m}41^{s}.58, δ = +22^{o}39′29′′.4 (equinox 2000.0; estimated uncertainty 0′′.5). This is just to the east of the x-ray error circle given on IAUC 7274 and ∼ 80′′ north of a 10th-mag star. The brightness of the optical transient is estimated to be R = 15.1, and it appears blue compared to other stars in the field. Images taken over a 3.3-hr span show < 0.05 mag variation in brightness. A faint object (mag ∼ 21) is visible on the Sky Survey within 1′′ of the position of the transient. The spectrum obtained with the FLWO 1.5-m telescope on Oct. 12.22 displays a strong blue continuum with weak emission lines. Hα, with an equivalent width of 0.54 nm, and a velocity width of 1200 km/s (FWHM), is seen. He II 468.6-nm is 3 times stronger than Hβ and blended with permitted C III and N III. The spectrum is consistent with an x-ray nova in outburst. A finder chart may be found at

http://cfa-www.harvard.edu/~peterg/gosox/j1859.jpg.

GM SAGITTARII R. M. Wagner, Ohio State University, communicates: “We obtained

a spectrum (range 320-680 nm, resolution 1.5 nm) of GM Sgr on Sept. 30.13 UT with the MDM Observatory 2.4-m Hiltner telescope (+ CCDS spectrograph). The spectrum exhibited relatively strong Balmer absorption lines visible to the series limit, in addition to interstellar absorption features at 578.0 and 589.0 (Na D) nm. Weaker absorption lines of Fe II might also be present. The Balmer lines appear to be much stronger than those observed in the later spectra obtained by Garcia et al. (IAUC 7271), indicating continuing evolution of the optical spectrum since the outburst. Equivalent widths were: Hα, 0.81; Hβ, 1.01; Hγ, 0.87; and Hδ, 0.85 nm. The FWHM of Hβ absorption was 1.7 nm. Convolution of the spectrum with broadband filters gives U − B = +1.8 and B − V = −0.2, although the U measurement might be affected by differential atmospheric refraction. The equivalent width of the 578.0-nm interstellar band was 0.06(1) nm, implying E(B − V) = 0.24 mag.”

1999 October 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green